# NeMo-Aligner: Scalable Toolkit for Efficient Model Alignment

### v1

@@ Abstract
Aligning Large Language Models (LLMs) with human values and preferences is essential for making them helpful and safe. However, building efficient tools to perform alignment can be challenging, especially for the largest and most competent LLMs which often contain tens or hundreds of billions of parameters. We create NeMo-Aligner, a toolkit for model alignment that can efficiently scale to using hundreds of GPUs for training. NeMo-Aligner comes with highly optimized and scalable implementations for major paradigms of model alignment such as: Reinforcement Learning from Human Feedback (RLHF), Direct Preference Optimization (DPO), SteerLM, and Self-Play Fine-Tuning (SPIN). Additionally, our toolkit supports running most of the alignment techniques in a Parameter Efficient Fine-Tuning (PEFT) setting. NeMo-Aligner is designed for extensibility, allowing support for other alignment techniques with minimal effort. It is open-sourced with Apache 2.0 License and we invite community contributions at https://github.com/NVIDIA/NeMo-Aligner.
@@ 1 Introduction
Despite the benefits of training models using feedback, these pipelines are notoriously challenging to get right (Lambert & Calandra, 2023; Zheng et al., 2023b), deterring widespread, productive adoption outside of select well-resourced organizations. For example, the popular Proximal Policy Optimization (PPO) variant of Reinforcement Learning from Human Feedback (RLHF) approach (Ouyang et al., 2022) requires running a complicated pipeline with four large language models interacting in a complex manner during training. Such alignment algorithms introduce new system challenges for efficient training that require re-thinking various aspects of the software stack including model scalability, coordination among models, and text generation (so-called ”rollout”) within the training loop.
There are existing open source tools for model alignment, most notably HuggingFace TRL (von Werra et al., 2020), CarperAI trlX (Havrilla et al., 2023) and Microsoft DeepSpeed-Chat (Yao et al., 2023). These tools provide an excellent starting point with respect to usability and feature set. However, with NeMo-Aligner we aim to vastly improve performance and scalability of PPO and other methods especially when it comes to the largest and most competent models such as Llama 2 70B (Touvron et al., 2023) and beyond.
NeMo-Aligner addresses scalability challenges by (I) building upon Megatron-LM (Shoeybi et al., 2020) with 3D (data, tensor, and pipeline)-parallelism training, (II) having a distributed approach to Proximal Policy Optimization (PPO) training in RLHF and (III) integrating PPO inference optimizations based on TensorRT-LLM (NVIDIA, 2023b) during rollout stage. Combined, these optimizations allow users to efficiently train the largest models over hundreds of GPUs reducing research iteration time.
@@ 2.1 Supervised Fine Tuning
Given a pre-trained (also referred to as ”base”) model, supervised fine-tuning (SFT) updates the base model’s parameters on prompts with expected responses, where the expected responses might come from expert human annotations (Köpf et al., 2023) or other language models (Ding et al., 2023). The model is trained to mimic the expected responses given prompts using the token-level cross-entropy loss. SFT is an important prerequisite step in Reinforcement Learning from Human Feedback (Ouyang et al., 2022) and Direct Preference Optimization (Rafailov et al., 2023) because without it, the base model is very unlikely to generate responses which follow user’s instructions. This step is also sometimes called behaviour cloning because the model is expected to mimic responses of a human or another model.
@@ 3 RLHF (PPO) Training
NeMo-Aligner is designed to support numerous alignment techniques efficiently at extremely large scales. It does so by building upon Megatron-LM (Shoeybi et al., 2020) and NeMo (Kuchaiev et al., 2019) to include features such as optimized kernels from Transformer Engine (NVIDIA, 2022), distributed fused adam optimizer and 3D parallelism support. NeMo-Aligner supports the entire RLHF pipeline as introduced by Ouyang et al. (2022) and described in Section 2.2. The training pipeline is separated into three distinct stages as illustrated in Figure 1: Supervised Fined Tuning, Reward Model Training, and Proximal Policy Optimization. The challenges with the pipeline efficiency come primarily from the Proximal Policy Optimization stage, and this section describes our approach to tackling these challenges, as summarized in Figure 2.
@@ 3.1 Distributed Approach to PPO training
All of these models can be extremely large (e.g. Llama 2 70B), so NeMo-Aligner takes a distributed approach to PPO training. We allow users to setup PyTriton (NVIDIA, 2022) servers and clients to communicate across the different models during PPO. These PyTriton servers make it possible to run the models on different compute clusters, removing the requirement of having both the critic and actor on the same compute allocation. Naively, four different servers (i.e. one for each model) would be launched. However, we note that the reference policy and PPO actor are the same model but with different weights. Therefore, we combine them into one job and offload the reference policy’s weights to CPU, swapping them with the actor’s weights for the reference policy inference step. We deploy the same strategy for the reward model and critic. All communications are done asynchronously, permitting pipelined critic inference/training with policy inference/training.
@@ 3.2 Optimizations for PPO rollout
Response generation during the rollout step is the main bottleneck during PPO training. The generation stage of the actor is composed of multiple forward passes, with one token generated per forward pass. Therefore, generation stage kernels are generally launch latency and memory bandwidth bound, meaning that directly reusing the compute optimized forward pass implementation of the training stage results in very poor performance.
To address these bottlenecks, we implement the generation stage using TensorRT-LLM (NVIDIA, 2023b), a high-performance LLM deployment framework. TensorRT-LLM integrates inference optimized kernels and automatic kernel fusion into a TensorRT based runtime to achieve better performance. At the start of RLHF, the model is passed to TensorRT-LLM, which takes the model and compiles it into a TensorRT engine; TensorRT-LLM loads the engine into its runtime and performs generation. The engine contains a copy of the model weights, while the runtime allocates the memory required by the KV-Cache and activations. Before and after generation, the engine and runtime are loaded and offloaded respectively. Because generation has lower memory requirements than training, we reshard the model to only use tensor parallelism.
On subsequent training steps, the engine must be synced with updated parameter weights from training. Because generation must wait until the weights are synced, we update the engine in-place using the TensorRT Refitter (NVIDIA, 2023c) instead of recompiling the engine which would incur a large overhead.
In addition, we observe that large discrepancies of generation time occur between the fastest and the slowest data parallel worker during generation due to the differences in response lengths. To mitigate this, we setup a worker pool to dynamically load balance among data parallel workers to give workers with shorter generations correspondingly more work.
@@ 3.3 Model Training Details and Quality
As a demonstration of practical large-scale RLHF training with NeMo-Aligner, we first perform SFT on a Llama 2 70B model using a mixture of public and proprietary data, then train a Llama 2 13B reward model using the Anthropic HH-RLHF dataset (Bai et al., 2022a), selecting the checkpoint with the lowest validation loss. Subsequently, we perform PPO using the same dataset.
For each PPO iteration, we set rollout batch size to $512$, optimizer batch size to $64$, KL Penalty ($\beta$) to $0.005$ and temperature to $1$. We use constant learning rate of $79e-$ 7 after $10$ warmup steps with the AdamW optimizer. Model was trained for $100$ rollout steps, based on validation performance on heldout validation set done every $4$ steps.
Following Jiang et al. (2023); Tunstall et al. (2023), we use MT-Bench (Zheng et al., 2023a) to evaluate the performance of the trained RLHF Model. The resulting model achieves a performance of 7.59 on MT-Bench, which is substantially higher than 6.86 achieved by Llama 2 70B Chat (Touvron et al., 2023). We have openly released this model alongside its training resources on Huggingface for reproducibility but avoid mentioning links here to maintain anonymity. For reproducibility, we release the SFT blend [^1] https://huggingface.co/datasets/nvidia/sft_datablend_v1, reward model [^2] https://huggingface.co/nvidia/NV-Llama2-13B-RLHF-RM, and RLHF-ed model [^3] https://huggingface.co/nvidia/NV-Llama2-70B-RLHF-Chat publicly.
@@ 3.4 Scalability
No. of compute nodes (Actor + Critic) 8 + 2 16 + 4 32 + 8
@@
Overall 187.6 (13.770) 111.4(8.904) 69.9 (2.583)
   Train 32.8 (0.161) 19.4 (0.040) 13.1 (0.038)
   Rollout 154.9 (13.917) 91.9 (8.912) 56.8 (2.554)
    - Response generation 110.2 (14.107) 61.9 (8.511) 33.5 (2.447)
    - Log-probs calculation 28.8 (0.006) 14.9 (0.021) 8.2 (0.009)
    - TensorRT Refit 12.1 (0.131) 11.8 (0.203) 11.8 (0.151)
    - Critic wait 0.1 (0.002) 0.1 (0.001) 0.1 (0.001)
Relative speed up (vs. 8 + 2 node setup) $\uparrow$
Overall 1x 1.68x 2.68x
   Train 1x 1.69x 2.50x
   Rollout 1x 1.69x 2.73x
    - Response generation 1x 1.78x 3.29x
    - Log-probs calculation 1x 1.93x 3.51x
Table 1: Effects of scaling training across different number of compute nodes for Llama 2 70B actor and Llama 2 13B critic on Rollout batch size of 512 and BF16 precision following Section 3.3. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step calculated based on mean of 5 steps after the first step, as the first step incurs substantial time for TRT-LLM Engine Building. Further training configuration details are in Table 5.
To demonstrate the scaling efficiency of NeMo-Aligner, we repeat identical training setups from Section 3.3, with i. 8 actor nodes + 2 critic nodes, ii. 16 actor nodes + 4 critic nodes and iii. 32 actor nodes + 8 critic nodes respectively. As shown in Table 1, overall time per step reduces correspondingly, achieving a 1.68x speed up between 8+2 nodes and 16+4 nodes as well as a 2.68x speed up when further scaling to 32+8 nodes. The speed up in overall time per step is contributed by speed ups in both the Train and Rollout stages, demonstrating the effective optimization that NeMo-Aligner has done for both stages.
The scaling of Train stage is sublinear (1.69x and 2.50x respectively when nodes double and quadruple) due to number of micro-batches per data parallel rank decreasing as node count increases. Because all pipeline stages must complete before the optimizer is called in pipeline parallel models, we incur an overhead to fill and drain the pipeline that is independent of the number of micro-batches (Shoeybi et al., 2020). Therefore, decreasing the number of micro-batches per data parallel rank increases the proportion of the train step spent in filling and draining the pipeline, where GPU utilization is poor.
Within the Rollout stage, the steps for Response generation and Log-probs calculation scales well with the number of nodes, achieving a 1.78x/3.29x and 1.93x/3.51x speed up when scaling to twice and four times the number of nodes respectively. This is because scaling up the number of actor nodes proportionally increases the number of data parallel workers for each step, which can evenly share the work.
On the other hand, the TensorRT Refit step within the Rollout stages takes a constant amount of time (11.8-12.1 seconds) irrespective of the number of compute nodes, as the inplace engine sync time is dominated by time spent reloading the engine. As a result, as the number of nodes increases from 8+2 to 32+8, the time taken by Tensor Refit as a proportion of the Rollout stage increases from 7.81% to 20.8%. Finally, async communications between the Actor and the Critic models result in the additional time taken to wait for the Critic model to be inconsequential (0.1 seconds), suggesting the effectiveness of having async non-blocking calls between actor and critic models in the PPO pipeline.
No. of compute nodes (Actor + Critic) 32 + 16 64 + 32
@@
Overall 147.0 (26.720) 90.1 (14.167)
   Train 21.2 (0.865) 15.1 (0.299)
   Rollout 125.8 (26.010) 75.0 (13.916)
    - Response generation 95.8 (26.274) 51.9 (14.182)
    - Log-probs calculation 14.8 (0.010) 8.1 (0.008)
    - TensorRT Refit 12.3 (0.281) 11.7 (0.162)
    - Critic wait 0.1 (0.002) 0.1 (0.014)
Relative speed up (vs. 32 + 16 node setup)$\uparrow$
Overall 1x 1.63x
   Train 1x 1.40x
   Rollout 1x 1.68x
    - Response generation 1x 1.85x
    - Log-probs calculation 1x 1.83x
Table 2: Effects of scaling training on 32 Actor node + 16 Critic node and 64 Actor node + 32 Critic node for Llama 2 70B actor and Llama 2 70B critic on Rollout batch size of 1024 and BF16 precision. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step calculated based on mean of 5 steps after the first step, as the first step incurs substantial additional time for TRT-LLM Engine Building. Further training configuration details are in Table 5.
System scalability also needs to be considered under the context of the problem requirements. The training setup in Section 3.3 has a 70B Llama 2 Actor, 13B Llama 2 Critic as well as a Rollout batch size of 512. Such a setup limits the effective demonstration of our system scaling beyond 32 + 8 nodes as there is not enough work to be meaningfully shared across more data parallel workers. Therefore, we modify the setup slightly to use a 70B Llama 2 Critic and a Rollout batch size of 1024 in Table 2 in order to measure the system performance when the requirements are higher. We use more nodes for the critic with the corresponding number of actor nodes (i.e. 32 + 16 instead of 32 + 8) because the critic is larger (70B vs. 13B) and hence require more compute to run effectively. Table 2 shows that the increased requirements of the training job allows it to meaningfully scale to 64 + 32 node (with 768 H100 GPUs total) for various stages within PPO.

### v2

@@ Abstract
Aligning Large Language Models (LLMs) with human values and preferences is essential for making them helpful and safe. However, building efficient tools to perform alignment can be challenging, especially for the largest and most competent LLMs which often contain tens or hundreds of billions of parameters. We create NeMo-Aligner, a toolkit for model alignment that can efficiently scale to a thousand GPUs for training the largest open-source LLMs such as Nemotron 4 340B and Llama 3.1 405B. NeMo-Aligner comes with highly optimized and scalable implementations for major paradigms of model alignment such as: Reinforcement Learning from Human Feedback (RLHF), Direct Preference Optimization (DPO), SteerLM, and Self-Play Fine-Tuning (SPIN). Additionally, our toolkit supports running most of the alignment techniques in a Parameter Efficient Fine-Tuning (PEFT) setting. NeMo-Aligner is designed for extensibility, allowing support for other alignment techniques with minimal effort. It is open-sourced with Apache 2.0 License and we invite community contributions at https://github.com/NVIDIA/NeMo-Aligner.
@@ 1 Introduction
Despite the benefits of training models using feedback, these pipelines are notoriously challenging to get right (Lambert & Calandra, 2023; Zheng et al., 2023b), deterring widespread, productive adoption outside of select well-resourced organizations. For example, the popular Proximal Policy Optimization (PPO) variant of Reinforcement Learning from Human Feedback (RLHF) approach (Ouyang et al., 2022) requires running a complicated pipeline with four large language models interacting in a complex manner during training. Such alignment algorithms introduce new system challenges for efficient training that require re-thinking various aspects of the software stack including model scalability, coordination among models, and text generation within the training loop.
There are existing open source tools for model alignment, most notably HuggingFace TRL (von Werra et al., 2020), CarperAI trlX (Havrilla et al., 2023) and Microsoft DeepSpeed-Chat (Yao et al., 2023). These tools provide an excellent starting point with respect to usability and feature set. However, with NeMo-Aligner we aim to vastly improve performance and scalability to more than a thousand GPUs, especially useful for aligning the largest and most competent models such as Nemotron 4 340B (Nvidia et al., 2024), Llama 3.1 405B (Dubey et al., 2024) and beyond.
NeMo-Aligner addresses scalability challenges by (I) building upon Megatron-LM (Shoeybi et al., 2020) with 3D (data, tensor, and pipeline)-parallelism training, (II) having a distributed approach to Proximal Policy Optimization (PPO) training in RLHF and (III) integrating PPO inference optimizations based on TensorRT-LLM (NVIDIA, 2023b) during rollout stage. Combined, these optimizations allow users to efficiently train the largest models over a thousand GPUs reducing research iteration time.
@@ 2.1 Supervised Fine Tuning
Given a pre-trained (also referred to as ”base”) model, supervised fine-tuning (SFT) updates the base model’s parameters on prompts with expected responses, where the expected responses might come from expert human annotations (Köpf et al., 2023) or other language models (Ding et al., 2023). The model is trained to mimic the expected responses given prompts using the token-level cross-entropy loss. SFT is an important prerequisite step in Reinforcement Learning from Human Feedback (Ouyang et al., 2022) and Direct Preference Optimization (Rafailov et al., 2023) because without it, the base model is very unlikely to generate responses which follow user’s instructions. This step is also sometimes called behavior cloning because the model is expected to mimic responses of a human or another model.
@@ 3 RLHF (PPO) Training
NeMo-Aligner is designed to support numerous alignment techniques efficiently at extremely large scales. It does so by building upon Megatron-LM (Shoeybi et al., 2020) and NeMo (Kuchaiev et al., 2019) to include features such as optimized kernels from Transformer Engine (NVIDIA, 2022), distributed fused adam optimizer and 3D parallelism support. NeMo-Aligner supports the entire RLHF pipeline as introduced by Ouyang et al. (2022) and described in Section 2.2. The training pipeline is separated into three distinct stages as illustrated in Figure 1: Supervised Fine-Tuning, Reward Model Training, and Proximal Policy Optimization. The challenges with the pipeline efficiency come primarily from the Proximal Policy Optimization stage, and this section describes our approach to tackling these challenges, as summarized in Figure 2.
@@ 3.1 Distributed Approach to PPO training
All of these models can be extremely large (e.g. Llama 3.1 405B), so NeMo-Aligner takes a distributed approach to PPO training. We allow users to setup PyTriton (NVIDIA, 2022) servers and clients to communicate across the different models during PPO. These PyTriton servers make it possible to run the models on different compute clusters, removing the requirement of having both the critic and actor on the same compute allocation. Naively, four different servers (i.e. one for each model) would be launched. However, we note that the reference policy and PPO actor are the same model but with different weights. Therefore, we combine them into one job and offload the reference policy’s weights to CPU, swapping them with the actor’s weights for the reference policy inference step. We deploy the same strategy for the reward model and critic. All communications are done asynchronously, permitting pipelined critic inference/training with policy inference/training.
@@ 3.2 Optimizations for PPO rollout
Response generation during the rollout step dominates the end to end time of PPO training. The generation stage of the actor is composed of multiple forward passes, with one token generated per forward pass. Therefore, generation stage kernels are generally launch latency and memory bandwidth bound, meaning that directly reusing the compute optimized forward pass implementation of the training stage results in very poor performance.
To address these bottlenecks, we implement the generation stage using TensorRT-LLM (NVIDIA, 2023b), a high-performance LLM deployment framework. TensorRT-LLM integrates inference optimized kernels and automatic kernel fusion into a TensorRT based runtime to achieve better performance. At the start of RLHF, the model is passed to TensorRT-LLM which compiles the model into a TensorRT engine; TensorRT-LLM loads the engine into its runtime and performs generation. The engine holds a copy of the model weights along with the runtime KV-cache and activations. Due to the cost of serializing the engine, we keep the engine in memory during training. As a result, we reduce the peak memory pressure by recomputing training stage activations in the backward pass. Furthermore, because generation has lower memory requirements than training, we reshard the model to only use tensor parallelism during inference if memory allows, removing overhead from inter-node communication when otherwise running with pipeline parallelism.
On subsequent training steps, the engine must be synced with updated parameter weights from the training stage. The engine is updated in-place using the TensorRT Refitter (NVIDIA, 2023c). We avoid recompiling the engine which would incur a large overhead as generation cannot begin until the weights are updated.
For large problem sizes, there may be discrepancies of generation time between the fastest and the slowest data parallel worker during generation due to the differences in response lengths. To mitigate this, we allow users to setup a worker pool to dynamically load balance among data parallel workers to give workers with shorter generations correspondingly more work.
@@ 3.3 Model Training Details and Quality
As a demonstration of practical large-scale RLHF training with NeMo-Aligner, we train Llama3 (Meta AI, 2024) 70B model using PPO as prescribed in Wang et al. (2024). The PPO model is trained with a rollout global batch size of 128, training global batch size of 128, constant learning rate of 1e-7, KL penalty of $0.003$, and using HelpSteer2 as a prompt source.
Following Wang et al. (2024); Meng et al. (2024), we use MT-Bench (Zheng et al., 2023a) with GPT-4-Turbo judge to evaluate the performance of the trained RLHF Model. The resulting model achieves a performance of 8.13 on MT-Bench, which is an improvement upon the strong starting SFT checkpoint with MT-Bench of 7.96. The Reward Model and RLHF-trained model are publicly released at https://huggingface.co/nvidia/Llama3-70B-SteerLM-RM and https://huggingface.co/nvidia/Llama3-70B-PPO-Chat respectively.
@@ 3.4 Scalability
No. of compute nodes (Actor + Critic) 8 + 4 16 + 8
@@
Overall 53.7 (2.223) 29.8 (1.459)
   Train 8.5 (0.356) 4.8 (0.118)
   Rollout 45.2 (1.957) 25.0 (1.41)
    - Response generation 35.8 (1.658) 17.5 (1.224)
    - Log-probs calculation 4.0 (0.573) 2.8 (0.275)
    - TensorRT Refit 3.1 (0.182) 2.3 (0.334)
    - Critic wait 0.01 (0.001) 0.01 (0.001)
Relative speed up (vs. 8 + 4 node setup) $\uparrow$
Overall 1x 1.80x
   Train 1x 1.77x
   Rollout 1x 1.81x
    - Response generation 1x 2.05x
    - Log-probs calculation 1x 1.43x
Table 1: Effects of scaling training across different number of compute nodes for Llama 3 70B actor and Llama 3 70B critic on rollout global batch size of 128 and BF16 precision following Section 3.3. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step calculated based on mean of 5 steps after the first step, as the first step incurs substantial time for TRT-LLM Engine Building. Further training configuration details are in Table 3.
To demonstrate the scaling efficiency of NeMo-Aligner, we repeat identical training setups from Section 3.3, with 8 actor nodes + 4 critic nodes and 16 actor nodes + 8 critic nodes. As shown in Table 1, overall time per step reduces correspondingly, achieving a 1.80x speed up between 8+4 nodes and 16+8 nodes. The speed up in overall time per step is contributed by speed ups in both the Train and Rollout stages, demonstrating the effective optimization that NeMo-Aligner has done for both stages.
The scaling of Train stage is sublinear due to number of micro-batches per data parallel rank decreasing as node count increases. Because all pipeline stages must complete before the optimizer is called in pipeline parallel models, we incur an overhead to fill and drain the pipeline that is independent of the number of micro-batches (Shoeybi et al., 2020). Therefore, decreasing the number of micro-batches per data parallel rank increases the proportion of the train step spent in filling and draining the pipeline, where GPU utilization is poor. A similar issue is apparent during the log prob calculation phase with scaling of 1.43x.
No. of compute nodes (Actor + Critic) 16 + 8 32 + 16 64 + 32
@@
Overall 190.4 (15.392) 106.8 (6.842) 56.9 (7.596)
   Train 38.8 (4.674) 22.2 (2.408) 14.1 (1.439)
   Rollout 151.5 (13.172) 84.6 (5.596) 42.7 (6.36)
    - Response generation 131.3 (12.551) 71.7 (5.093) 29.9 (6.456)
    - Log-probs calculation 15.5 (3.055) 8.5 (1.753) 6.2 (1.177)
    - TensorRT Refit 2.2 (0.408) 2.2 (0.014) 3.5 (0.063)
    - Critic wait 0.02 (0) 0.03 (0.001) 1.7 (2.045)
Relative speed up (vs. 16 + 8 node setup)$\uparrow$
Overall 1x 1.78x 3.35x
   Train 1x 1.75x 2.75x
   Rollout 1x 1.79x 3.55x
    - Response generation 1x 1.83x 4.39x
    - Log-probs calculation 1x 1.82x 2.50x
Table 2: Effects of scaling training for Llama 3 70B actor and Llama 3 70B critic on rollout global batch size of 1024 and BF16 precision. Nodes are 8*H100-80GB-SXM connected with intra-node NVLink (NVIDIA, 2023a) and inter-node Infiniband (NVIDIA, 2024) interconnects. Time per step calculated based on mean of 5 steps after the first step, as the first step incurs substantial additional time for TRT-LLM Engine Building. Our distributed optimizer feature enables NeMo-Aligner to support double the rollout microbatch size from 8 to 16 in the 64 + 32 node configuration. This speeds up the response generation time significantly. Further configuration details are in Table 3.
Generation time scales well with the number of nodes, achieving a near-linear speedup of 2.05x. This is because scaling up the number of actor nodes proportionally increases the number of data parallel workers for each step, which can evenly share the intense work of generation. However, the time spent fitting weights into the TensorRT engine remains relatively constant and therefore does not scale well with node count. Finally, async communications between the Actor and the Critic models result in the additional time taken to wait for the Critic model to be inconsequential (0.01 seconds), suggesting the effectiveness of having async non-blocking calls between actor and critic models in the PPO pipeline.
Model No. of compute nodes Tensor Parallel Pipeline Parallel Data Parallel Rollout Batch Size Micro Global 70B + 70B 8 + 4 8 + 8 8 + 4 1 + 1 8 128 70B + 70B 16 + 8 8 + 8 8 + 4 2 + 2 8 128 / 1024 70B + 70B 32 + 16 8 + 8 8 + 4 4 + 4 8 1024 70B + 70B 64 + 32 8 + 8 8 + 4 8 + 8 16 1024 405B + 405B 84 + 42 8 + 8 21 + 21 4 + 2 16 128
Table 3: Parallelism settings for scaling experiments on Llama3 models. The node counts and parallelism configurations are denoted as actor + critic.
System scalability also needs to be considered under the context of the problem requirements. The training setup in Section 3.3 has a 70B Llama 3 Actor, 70B Llama 3 Critic as well as a rollout global batch size of 128. Such a setup limits the effective demonstration of our system scaling beyond 16 + 8 nodes as there is not enough work to be meaningfully shared across more data parallel workers. Therefore, we modify the setup slightly to use a rollout global batch size of 1024 in Table 2 in order to measure the system performance when the requirements are higher. Table 2 shows that the increased requirements of the training job allows it to meaningfully scale to 64 + 32 nodes (with 768 H100 GPUs total) for various stages within PPO.